OpenStudy (babbs12):

Use the mid-point rule with n = 2 to approximate the area of the region bounded by y equals the cube root of the quantity 16 minus x cubed y = x, and x = 0.

OpenStudy (babbs12):

\[\sqrt[3]{16-x^3}\]

OpenStudy (mathmate):

First step: find the end points of the region. Can you do that?

OpenStudy (babbs12):

the limits?

OpenStudy (mathmate):

yes! the limits of the integration, which correspond to the end points of the region, where \(\Large \sqrt[3]{16-x^3}\) and \(\Large y=x\) meet.

OpenStudy (babbs12):

the point where they meet is 2,2 according to the graph

OpenStudy (mathmate):

Excellent, so you've drawn a graph of the two functions! So now you know the limits of integration!

OpenStudy (babbs12):

the limits are 0 and cbrt(16-x^3)

OpenStudy (mathmate):

That's the y-limits. To do the mid point rule, you split the region into vertical strips, n=2 means two of them. So the limits along x-axis are 0 and 2.

OpenStudy (babbs12):

okay

OpenStudy (mathmate):

Do you know how the mid-point rule works?

OpenStudy (babbs12):

its the mid point of the rectangle used in the reimann sum right?

OpenStudy (babbs12):

middle/height

OpenStudy (mathmate):

Exactly. 

OpenStudy (mathmate):

Can you give the answer in terms of the diagram?

OpenStudy (babbs12):

so the mid points would be .5 and 1.5

OpenStudy (mathmate):

You are approximating the area of the region with two rectangles, each of height y1 and y2. Yes the mid-points are 0.5 and 1.5 (where the function is evaluated)

OpenStudy (babbs12):

so the function is cbrt(16-x^2) - x ???

OpenStudy (mathmate):

 The sum of the area of the two rectangles is your answer. Yes, that's the height of the rectangles, at x=0.5, and x=1.5. Great!